OpenStudy (anonymous):

i need help writing equation based on a word proble. I have tried three times now and have been told each time my variables are not the same when i write them help me pls..the problem and equations i wrote r below.

OpenStudy (anonymous):

hi i am so happy to see u

Directrix (directrix):

@jannine245 I understand the problem situation but what is the question as stated in your text? That is, what are we supposed to write or to solve or to do or to show as an answer?

OpenStudy (anonymous):

i need to write the equations to solve the problems

OpenStudy (anonymous):

then find out when the equations owuld be equal

OpenStudy (anonymous):

but first is the wiritng of the equations i am tld my variables r nto all defined hte same

Directrix (directrix):

What about the 37 hours a week that Mrs. Lewis works? Are we to find out which is the cheaper sitter service for 37 hours? Is that the question?

OpenStudy (anonymous):

yes i put that there but not sure if i need it..let me get u the original directions

Directrix (directrix):

ok

OpenStudy (anonymous):

got it

Directrix (directrix):

Is the 37 given in the problem statement? > Mrs. Lewis will be working about 37 hours a week so she needs to see which would be the best choice for her son and her wallet.

OpenStudy (anonymous):

no i made htis problem myself

Directrix (directrix):

Here is what is wrong with the variable x not being the same: 1.Kiddie Castle: Y= $224 + $6.00(x - 35) Here, x > 35 because the extra money kicks in when Joey stays at KC more than 35 hours. 2. Cultural Au Pair: y = 7.00*x Here, x represents the number of hours the AP is at Joey's house. So, x could be 30. No problem with y = 7.00*x At the KC, there is a problem: Y= $224 + $6.00(x - 35) Y = 224 + 6.00* (30 - 35) = 224 + 6.00* (-5) so Joey's bill goes down $30 although he stays at KC five more hours. Do you see the variable problem? We can fix that by changing the charges at Kiddy Castle.

Directrix (directrix):

I think it is best to have KC charge a rate that does not change after a certain number of hours.

Directrix (directrix):

KC: charges a flat rate of $100 per week plus $6 per hour ACP: charges $7 per hour Is that okay?

OpenStudy (anonymous):

i tried that and was told to it is wrong does nto meet the requirements

Directrix (directrix):

What are the original requirements of the problem as stated in the text?

Directrix (directrix):

I am wondering about the 37 hours Mrs Lewis plans to work per week. Did the teacher say anything about that not being part of the solution?

Directrix (directrix):

The 37 or some number greater than 35 should be there to show the cost difference. It's that x has to represent the same thing in both equations.

OpenStudy (anonymous):

yes and that is where i am having the problem

OpenStudy (anonymous):

i am just working on this ti is my last one to do nad is driving me crazy i will close then if it is easier

OpenStudy (anonymous):

ok

Directrix (directrix):

Kiddie Castle: Y= $224 + $6.00(x - 35) Y = Cultural Au Pair: y = 7.00*x I moved this down here so that I could look at it. I understand what the problem requires.

OpenStudy (anonymous):

no problem but wouldnt it be three equations

Directrix (directrix):

Oh, that's just the facts of the problem. The equations are not correct with the x problem.

Directrix (directrix):

Kiddie Castle: Y= $224 + $6.00(x - 35), if x ≥ 35 Cultural Au Pair: Y = 7.00*x + 0*(x - 35), if x ≥ 35 This is just an idea. I don't think it will fly.

OpenStudy (anonymous):

thank you sooooooo much you have no idea

Directrix (directrix):

I am graphing now. Hold on.

Directrix (directrix):

The teacher may say the piecewise idea is too advanced or whatever but not really. One thing, the variable x represents the same thing in both equations.

Directrix (directrix):

Let me see what happens when I graph.

Directrix (directrix):

I don't think so. Try it and see what happens.

OpenStudy (anonymous):

it would be 14 way to small

Directrix (directrix):

32 hours is the break even point for CAP and KC. Anything beyond 32 hours means KC costs more.

OpenStudy (anonymous):

how did u find that

OpenStudy (anonymous):

need to do it algebracailly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i see that

Directrix (directrix):

Does the graph make sense?

OpenStudy (anonymous):

yes see how it goes with the hours and all i get that

OpenStudy (anonymous):

i see what you did but if i am using the two equations y=7x and y=224 +6(x-35) then how do i do this or set them equal to ero

Directrix (directrix):

I know you want or need just two equations but at the moment, I don't see how to get that with the constraints of the problem.

OpenStudy (anonymous):

i was not going to do 0 at l

Directrix (directrix):

I messed it up. That is why I mentioned it. ----------------- Do you have a problem you went over in class that meets the requirements of this assignment? I realize that you are a little unsure about this piece-wise stuff. I understand that because I don't know how much of that the class has studied.

OpenStudy (anonymous):

no i dont that is why i am so lost

Directrix (directrix):

If you were asked to explain the problem we just did, do you understand it well enough to explain it?

OpenStudy (anonymous):

i feel preety comfortable with it thanks to u

Directrix (directrix):

If one equation has to have overtime in it and the other one cannot (makes it be a horizontal line), I don't see how to get around the piecewise definition or 2-part definition. If you want more algebra, we can increase the KC flat rate so that KC will intersect at, say 40 hours, and you can have more algebra to show. It would be the same problem, piecewise, but with slightly different numbers.

OpenStudy (anonymous):

noleave it this way

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

ty again

Directrix (directrix):

Are you saying that 35 hours is the break even point?

OpenStudy (anonymous):

yes i hope this will work it will be my third time doing

Directrix (directrix):

I believe it will work. You can certainly defend it based on the equation. I see no way to not split up one of the equations when the other equation has to be flat rate.

OpenStudy (anonymous):

i hope it works..hmm how should i word how to find when the two equations r equal if i have these 2 equations

Directrix (directrix):

That sounds good. At this point, we do not know that the break even point is 32 so we are testing to see if it is. If it is not, then we go to the third equation. We will not need to do that but I would mention it.

OpenStudy (anonymous):

i am going ot try now to do itif nto for sure by tomorrow i will have it

Directrix (directrix):

Cool graph. Let me read the problem statement above. Just a minute.

OpenStudy (anonymous):

oh no i have it wrong u have to subtract 35 from the number of hours

Directrix (directrix):

I think it is just written wrong in the description. Edit it.

OpenStudy (anonymous):

i hope they dont take points off for it

Directrix (directrix):

The graph description is okay.

OpenStudy (anonymous):

i did b4 i was trying ot get u before

OpenStudy (anonymous):

i will try i guess..thank you for all your help i hope i pass this time

Directrix (directrix):

I believe you will do more than pass. But, if anything goes wrong, then take careful notes on what the teacher says is not correct , and we will keep working until we get it right. Look, I think it would be a good idea to delete your project upload and the problem statement from this thread. So, I'll look to hear from you tomorrow or whenever the teacher gives you feedback. Let me know. Okay? Keep a positive outlook.

OpenStudy (anonymous):

ty i was going ot do that delete it

Directrix (directrix):

>i did b4 i was trying ot get u before I'm sorry I missed you earlier tonight.

OpenStudy (anonymous):

its ok wishme luck

Directrix (directrix):

@jannine245 How did this project turn out?